There are expanded isothermally 3 lb/sec of NH, from saturated vapor to the state of 10 psia, 100°F. Determine (a) Q, (b)W (nonflow), (c) W (steady flow with ∆K=0), (d) ∆G, (e) the unavailable part of Q for t= 60°F. See Item B 33.

1
Expert's answer
2021-12-02T10:07:59-0500

"W=pV=\frac{\dot mRT}M=207~\frac{kJ}s,"

"\dot V=\frac {\dot mRT}{pM}=1.58~\frac{m^3}s."